OpenStudy (lena772):

A business has $11,080 to spend on new laptops and tablet computers for its salespeople. The laptops cost $515 each. The tablets cost $285 each. The business wants each salesperson to have either a laptop or a tablet. There are 30 salespeople. How many of each type of computer should the business buy? 9 laptops and 21 tablets 11 laptops and 19 tablets 19 laptops and 11 tablets 21 laptops and 9 tablets

hero (hero):

Let x = the salespeople with laptops y = the salespeople with tablets Then The total people is represented by x + y = 30 The total spent is represented by 515x + 285y = 11080

hero (hero):

Therefore, solve the following system of equations: x + y = 30 515x + 285y = 11080

OpenStudy (lena772):

(x+y=30) 285x+285y=8550 515x+285y=11080 -230x=-2530 x=11

OpenStudy (lena772):

11+y=30 y=19

hero (hero):

Great job.